Adria Lopez, owner of Success Systems, realizes that she needs to begin accounting for bad debts expense. Assume that Success Systems has total revenues of $43,853 during the first three months of 2017, and that the Accounts Receivable balance on March 31, 2017, is $22,720. Required: 1a. Prepare the adjusting entry needed for Success Systems to recognize bad debts expense, which are estimated to be 1% of total revenues on March 31, 2017 (assume a zero unadjusted balance in the Allowance for Doubtful Accounts at March 31). 1b. Prepare the adjusting entry needed for Success Systems to recognize bad debts expense, which are estimated to be 2% of accounts receivable on March 31, 2017 (assume a zero unadjusted balance in the Allowance for Doubtful Accounts at March 31) 2. Assume that Success Systems' Accounts Receivable balance at June 30, 2017, is $20,250 and that one account of $100 has been written off against the Allowance for Doubtful Accounts since March 31, 2017. If Adria Lopez uses the method prescribed in Part 1b, what adjusting journal entry must be made to recognize bad debts expense on June 30, 2017?
Answer #1
1a
Bad debts expense 439 =43853*1%
        Allowance for Doubtful Accounts 439
b
Bad debts expense 454 =22720*2%
        Allowance for Doubtful Accounts 454
2
Bad debts expense 51 =(20250*2%)-(454-100)
        Allowance for Doubtful Accounts 51